OpenStudy (anonymous):

Parametrize the intersection between the sphere \[ x^2+y^2+z^2=1 \]And the plane \[ x+y+z=0 \] Is there any good trick for this? My intuition is to start with \[ (\cos(t),\sin(t),0)\quad t\in [0,2\pi] \]With a normal vector \(<0,0,1>\) and then rotate the parametrization to align with the normal vector of the plane \(<1,1,1>\)

OpenStudy (anonymous):

Another idea I have it to project the intersection onto the \(xy\) plane and then after parametrizing \(x\) and \(y\), using the equations to parametrize \(z\).

OpenStudy (anonymous):

It's been a while since I've parametrized intersections. I solved for z in the second equation and substituted, getting: \[2(x^{2} + xy + y ^{2}) = 1\] This equation seems awfully familiar to me but I can't remember how to parametrize.

OpenStudy (anonymous):

I have no idea how to parametrize that.

OpenStudy (anonymous):

Anyone got any ideas?

OpenStudy (anonymous):

The rotation matrix for \(45^\circ\) is \[ \begin{bmatrix} \cos(45^\circ) & -\sin(45^\circ) \\ \sin(45^\circ) & \cos(45^\circ) \end{bmatrix} = \begin{bmatrix} \sqrt{1/2} & -\sqrt{1/2} \\ \sqrt{1/2} & \sqrt{1/2} \end{bmatrix} \]

OpenStudy (anonymous):

I think your rotation strategy is a good idea, I'm just not familiar with the implementation.

OpenStudy (anonymous):

 Something like this. Transformations are not tough to do on parametrizations.

OpenStudy (anonymous):

 We have this sphere intersected by a plane so I know it's some sort of rotated circle.